OpenStudy (anonymous):

You (safely) bungee jump from a 200-feet tall bridge in your town. Your distance above the water's surface depends on the time since you jumped. Sketch a reasonable graph.

OpenStudy (anonymous):

have you found out how to do it?

OpenStudy (anonymous):

@compboi23 no i havent

OpenStudy (ybarrap):

y=yi+vi*t+1/2*a*t ^{2} yi = initial position = 200 a = -32.2 ft/s^2 t = time vi = 0 (initial velocity) so position as a function of time will be a parabola starting at 200 pointing downward and hitting the x-axis a y=0

OpenStudy (ybarrap):



OpenStudy (ybarrap):